Some remaining questions that should be addressed:
1. IF the HVs are logged in the data stream at the start (end?) of each run would someone please extract them for runs 2277 and 2278 and look for a change in values for the NA-P1 bar 5 PMTs. Hell, do a diff on all 'meas HVs' and log any discrepancies -- nothing should have changed.
2. Nothing has been done to address the low ADC spectrum on the right PMT
for bar 5 (the set, meas, demand values in the GUI were all
self-consistent). ** Should the HV be bumped up? **